Sample means from a infinite population

A population has mean of 60 and a standard deviation of 5. A random sample of 16 

measurements is drawn this population. Describe the sampling distribution of the 

sample mean by computing its mean and standard deviation


1
Expert's answer
2022-03-14T17:42:29-0400

"\mu_{\bar X}=\mu=60."

"\sigma_{\bar X}=\frac{\sigma}{\sqrt{n}}=\frac{5}{\sqrt{16}}=1.25."